The slope of a line is -3/4 and it passes through the point (-7,9).Write the equation of the line in slope intercept form.
Answer by lenny460(1073)

The slope of a line is -3/4 and it passes through the point (-7,9).Write the equation of the line in slope intercept form.

We use the formula:

y - y1 = m(x - x1)
The Slope (m) = -3/4
The point is: (-7,9)
y -9 = -3/4(x - (-7)
(x - (-7) = x + 7
Therefore:
y -9 = -3/4(x + 7)
-3/4 * x = -3x/4
-3/4 * 7 = -21/4 = -5 1/4

y - 9 = -3x/4 - 5 1/4
Add 9 to each side of the equation.
y - 9 + 9 = -3x/4 - 5 1/4 + 9
Combine like terms:
-9 + 9 = 0
-5 1/4 + 9 = 3 3/4
Therefore the equation of the line is:

y = -3x/4 + 3 3/4 or
y = -3x/4 + 15/4
Using the above when you insert x = -7
you will get y = 9



Lennox Obuong
Algebra Student
Email: obuong3@aol.com